# MixGCN: Scalable GCN Training by Mixture of Parallelism and Mixture of Accelerators

### v2

@@ 1. Introduction
Despite GCNs’ great potential, designing scalable GCN training systems is particularly challenging and still under-explored, due to the associated giant feature tensors. For example, ogbn-papers100M (hu2020open, ), a popular dataset for GCN research, contains more than 100 million nodes, requiring 124GB for storing merely the features and labels, let alone the storage requirement for storing the intermediate features to support backward propagation, which cannot be fit in a single modern accelerator. As such, to handle large-graph training, many recent works follow the direction of partition parallelism (ma2019neugraph, ; zhu2019aligraph, ; jia2020improving, ; zheng2020distdgl, ; wang2021flexgraph, ; tripathy2020reducing, ; md2021distgnn, ; fey2021gnnautoscale, ; thorpe2021dorylus, ; gandhi2021p3, ; wang2020gnnadvisor, ; wan2022bns, ; wan2022pipegcn, ; ramezani2021learn, ; peng2022sancus, ; wang2022neutronstar, ; zhang2024sylvie, ; huang2024wisegraph, ; wan2023adaptive, ), as depicted in Figure 1b. The key idea is to separate a giant graph into multiple partitions, and assign each partition to one single accelerator. This straightforward approach, however, only distributes the storage of feature tensors, while incurring a significant memory and communication overhead for duplicating the remote neighbors from other accelerators (i.e., the red nodes in Figure 1b) (wan2022bns, ). This has restricted the scalability of GCN training due to the scaled-out number of remote neighbors. Furthermore, as we will show in Section 3.1.1, balancing workload via partition parallelism is NP-Hard, leading to a nontrivial synchronization overhead.
@@ 2.2. Partition Parallelism for GCN Training
To improve the scalability of GCN training, many recent works follow the paradigm of partition parallelism (zhang2023survey, ), which is depicted in Figure 1b. They either develop a scheduling algorithm towards balanced workload or optimized communication (ma2019neugraph, ; zhu2019aligraph, ; jia2020improving, ; wang2021flexgraph, ; md2021distgnn, ; wang2022neutronstar, ; huang2024wisegraph, ) or adjust training algorithm to reduce or hide its communication overhead (thorpe2021dorylus, ; wan2022pipegcn, ; ramezani2021learn, ; wan2022bns, ; peng2022sancus, ; wan2023adaptive, ; zhang2024sylvie, ). However, as we will point out in Section 3.1.1, although partition-parallel training distributes the storage of giant feature tensors, it suffers from scaled-out memory overhead and communication volume due to the duplicated remote neighbors (i.e., the red nodes in Figure 1b). Furthermore, since a real-world graph is often highly irregular, finding a balanced-workload partition is NP-hard.

### v3

@@ 1. Introduction
Despite GCNs’ great potential, designing scalable GCN training systems is particularly challenging and still under-explored, due to the associated giant feature tensors. For example, ogbn-papers100M (hu2020open, ), a popular dataset for GCN research, contains more than 100 million nodes, requiring 124GB for storing merely the features and labels, let alone the storage requirement for storing the intermediate features to support backward propagation, which cannot be fit in a single modern accelerator. As such, to handle large-graph training, many recent works follow the direction of partition parallelism (ma2019neugraph, ; zhu2019aligraph, ; jia2020improving, ; zheng2020distdgl, ; wang2021flexgraph, ; tripathy2020reducing, ; md2021distgnn, ; fey2021gnnautoscale, ; thorpe2021dorylus, ; gandhi2021p3, ; wang2020gnnadvisor, ; wan2022bns, ; wan2022pipegcn, ; ramezani2021learn, ; peng2022sancus, ; wang2022neutronstar, ; chai2022distributed, ; bai2023staleness, ; zhang2024sylvie, ; huang2024wisegraph, ; wan2023adaptive, ), as depicted in Figure 1b. The key idea is to separate a giant graph into multiple partitions, and assign each partition to one single accelerator. This straightforward approach, however, only distributes the storage of feature tensors, while incurring a significant memory and communication overhead for duplicating the remote neighbors from other accelerators (i.e., the red nodes in Figure 1b) (wan2022bns, ). This has restricted the scalability of GCN training due to the scaled-out number of remote neighbors. Furthermore, as we will show in Section 3.1.1, balancing workload via partition parallelism is NP-Hard, leading to a nontrivial synchronization overhead.
@@ 2.2. Partition Parallelism for GCN Training
To improve the scalability of GCN training, many recent works follow the paradigm of partition parallelism (zhang2023survey, ), which is depicted in Figure 1b. They either develop a scheduling algorithm towards balanced workload or optimized communication (ma2019neugraph, ; zhu2019aligraph, ; jia2020improving, ; wang2021flexgraph, ; md2021distgnn, ; wang2022neutronstar, ; huang2024wisegraph, ) or adjust training algorithm to reduce or hide its communication overhead (thorpe2021dorylus, ; wan2022pipegcn, ; ramezani2021learn, ; wan2022bns, ; peng2022sancus, ; chai2022distributed, ; bai2023staleness, ; wan2023adaptive, ; zhang2024sylvie, ). However, as we will point out in Section 3.1.1, although partition-parallel training distributes the storage of giant feature tensors, it suffers from scaled-out memory overhead and communication volume due to the duplicated remote neighbors (i.e., the red nodes in Figure 1b). Furthermore, since a real-world graph is often highly irregular, finding a balanced-workload partition is NP-hard.
@@ 6. Conclusion
Appendix A Proof of Proposition 3.1
We reduce a well-known NP-Hard problem identical-machines scheduling [56] to balancing workload of partition parallelism to conclude the proof.
Proof.
Suppose that we have a total of $m$ workers. Partition parallelism distributes the computation of all nodes across $m$ workers. Denote the assigned subgraph of worker $i$ as $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, then the total floating point operations (FLOPs) for the $l$-th layer is:
$\sum_{v\in\mathcal{V}_{i}}\left(deg(v)d^{(l)}+d^{(l)}d^{(l+1)}\right)$
where $deg(v)$ represents the degree of node $v$ and $d^{(l)}$ denotes the dimension of the $l$-th layer’s input. The first term $deg(v)d^{(l)}$ is the operations for aggregating the features of $v$’s neighbors, and we need extra $d^{(l)}d^{(l+1)}$ operations for updating its feature.
We now define $C_{v}\triangleq d^{(l)}(d^{(l+1)}+deg(v))$. For balancing the workload of partition parallelism, the partitioned node set $\{\mathcal{V}_{1},\cdots,\mathcal{V}_{m}\}$ should minimize the following objective:
$\max_{i\in[m]}\sum_{v\in\mathcal{V}_{i}}C_{v}$
which is a standard identical-machines scheduling problem and well-known to be NP-Hard [56]. Hence, balancing the computation workload of partition parallelism is NP-Hard. ∎